OpenStudy (sarahc):

Given: x – (5 – 3x + 2) = 15 + 3x Prove: x = 22 x – (5 – 3x + 2) = 15 + 3x Given x – (5 + 2– 3x) = 15 + 3x x – (7– 3x) = 15 + 3x x – 7 + 3x = 15 + 3x x + 3x – 7 = 15 + 3x Commutative Property of Addition 4x – 7 = 15 + 3x Combine Like Terms x – 7 = 15 Subtraction Property of Equality x = 22 Addition Property of Equality

OpenStudy (sarahc):

What are the missing justification steps, in order? A. Commutative Property of Addition, Combine Like Terms, Distributive Property B. Associative Property of Addition, Distributive Property, Multiplication Property of Equality C. Associative Property of Addition, Commutative Property of Addition, Distributive Property D. Associative Property of Addition, Distributive Property, Addition Property of Equality

OpenStudy (sarahc):

@jim_thompson5910 a? maybe?

OpenStudy (lgbasallote):

perhaps it's wise to point \(\textbf{*where*}\) the missing steps are?

OpenStudy (sarahc):

what?

OpenStudy (lgbasallote):

where are the missing steps? there are like a lot of unnamed steps there...where are you supposed to start?

jimthompson5910 (jim_thompson5910):

you are correct, the answer is indeed choice A

OpenStudy (sarahc):

okay thanks again, I always feel like I'm getting these wrong. too many words..hehe

jimthompson5910 (jim_thompson5910):

you're doing fine, just take it one step at a time

OpenStudy (sarahc):

Given: 4(x + 1) = 10x – 8 Prove: x = 2 Step Mathematical Statement Justification STEP 0: 4(x + 1) = 10x – 8 Given STEP 1: 4x + 4 = 10x – 8 Distributive Property STEP 2: 4 = 6x – 8 Subtraction Property of Equality STEP 3: 12 = 6x Addition Property of Equality STEP 4: 6x = 12 Symmetric Property of Equality STEP 5: x = 2 Associative Property of Multiplication

OpenStudy (sarahc):

Which step of justification is incorrect, and what should the justification for that step be to solve the equation? Step 5; Division Property of Equality Step 1; Commutative Property of Addition Step 3; Subtraction Property of Equality Step 2; Addition Property of Equality

OpenStudy (sarahc):

C? :o

jimthompson5910 (jim_thompson5910):

no, step 3 is just fine...so the answer isn't C

OpenStudy (sarahc):

oh :( step 2 ?

jimthompson5910 (jim_thompson5910):

step 2 is just fine as well

jimthompson5910 (jim_thompson5910):

because in that step, we're subtracting 4x from both sides, which is perfectly valid

jimthompson5910 (jim_thompson5910):

take a look at step 5 and look up the definition of the Associative Property of Multiplication

OpenStudy (sarahc):

ohhhh its when 3 or more numbers are multiplied!

jimthompson5910 (jim_thompson5910):

yes, but we're dividing both sides by 6 to isolate x...so we're actually using the Division Property of Equality

jimthompson5910 (jim_thompson5910):

So step 5 is correct, but the reason/justification is incorrect

OpenStudy (sarahc):

Okay I got it :) I swear i should start paying you....Lol!

jimthompson5910 (jim_thompson5910):

lol it's fine, although I do appreciate the thought

jimthompson5910 (jim_thompson5910):

but I'm glad you're understanding it now

OpenStudy (sarahc):

haha okay :) & me tooo! I think i got the last two..

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (sarahc):

Wait last one I swear. i just wanna check my answer :))

OpenStudy (sarahc):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (sarahc):

Subtraction property of equality??

jimthompson5910 (jim_thompson5910):

You are correct. The subtraction property of equality allows you to subtract terms from both sides.

OpenStudy (sarahc):

yay, thanks. &good night! :)

jimthompson5910 (jim_thompson5910):

you're welcome, good night